# GREAT: Geometry-Intention Collaborative Inference for Open-Vocabulary 3D Object Affordance Grounding

### v1

@@ Abstract
Open-Vocabulary 3D object affordance grounding aims to anticipate “action possibilities” regions on 3D objects with arbitrary instructions, which is crucial for robots to generically perceive real scenarios and respond to operational changes. Existing methods focus on combining images or languages that depict interactions with 3D geometries to introduce external interaction priors. However, they are still vulnerable to a limited semantic space by failing to leverage implied invariant geometries and potential interaction intentions. Normally, humans address complex tasks through multi-step reasoning and respond to diverse situations by leveraging associative and analogical thinking. In light of this, we propose GREAT (GeometRy-intEntion collAboraTive inference) for Open-Vocabulary 3D Object Affordance Grounding, a novel framework that mines the object invariant geometry attributes and performs analogically reason in potential interaction scenarios to form affordance knowledge, fully combining the knowledge with both geometries and visual contents to ground 3D object affordance. Besides, we introduce the Point Image Affordance Dataset v2 (PIADv2), the largest 3D object affordance dataset at present to support the task. Extensive experiments demonstrate the effectiveness and superiority of GREAT. Code and dataset are available at project.
@@ 1 Introduction
Figure 1: Differernce and Motivation. (a) object affordance grounding on seen setting. (b) Open-Vocabulary Affordance Grounding (OVAG) with previous paradigms. (c) observing interaction images, people engage in brainstorming through memory representations, drawing on prior interaction experiences to perform analogical reasoning and infer appropriate actions. (d) OVAG with our geometry-intention collaborative inference with chain-of-thought, step-by-step identifies the interaction part, extracts geometric attributes, reasons about corresponding interaction and brainstorms underlying interaction intentions, jointly grounding the 3D object affordance.
@@
Open-Vocabulary 3D object affordance grounding aims to locate “action possibilities” on objects [14, 8], both for seen and unseen scenarios, identifying specific regions on objects that support certain interactions. This bridges visual perception and physical manipulation for embodied agents and possesses bountiful application scenarios, e.g. robot manipulation [12, 20, 41], scene understanding [29, 11, 15], action anticipation [63, 64, 36] and imitation learning [25, 13].
@@
Recently, most existing methods [6, 18, 2] establish explicit mappings between semantic affordance categories and geometric structures, restricted to predefined seen categories and fail to ground object affordance out of the training categories. Thus, some studies [56, 27, 42, 50, 58] explore grounding object affordance through additional instructions, encompassing combining images or languages that depict interactions with 3D geometries to introduce external interaction priors, and mitigate the generalization gap lead by affordance diversity. Despite their remarkable progress, they are still vulnerable to a limited semantic space by failing to leverage implied invariant geometries among objects with the same affordance, as well as potential correlations among distinct interactions of the same object. As shown in Fig. 1 (a), current paradigms ground the 3D object affordance by aligning the object’s geometric features and instruction modalities, working well in the seen partition. However, such a paradigm relies excessively on the data fed into the model, when testing unseen affordance like “pour” that does not appear in the training set, the model endeavors to categorize it as the seen category “grasp”, shown in Fig. 1 (b).
@@
In cognitive science, studies [7, 17] have shown that humans tackle complex tasks through multi-step reasoning and respond to diverse situations by employing associative and analogical thinking. As demonstrated in Fig. 1 (c), when observing a water pouring scenario, humans employ multi-step reasoning to infer multiple potential interaction procedures and identify geometric properties of objects, indexing relevant knowledge from their brains. Analogous to this procedure, we leverage Multi-modal Large Language Models (MLLMs) [4, 24, 37, 49] to simulate prior knowledge, encompassing visual, linguistic, and other modalities, lifting their superior reasoning and generalization abilities to visual tasks [26, 43]. However, the dynamic and diverse nature of affordances makes it challenging to solely derive complex reasoning outcomes from MLLMs, to address this, a chain-of-thought (CoT) inference strategy that mirrors human reasoning processes is designed (Fig. 1 (d)). The strategy step-by-step identifies the interaction primitives, extracts geometric attributes, and reasons about the corresponding interaction actions and intentions, which enables the model to analyze interaction images from multiple perspectives, thereby eliminating the hallucination and ambiguities raised by MLLMs when reasoning interactions.
@@
To achieve this, we present the GREAT, a novel framework that excavates implied affordance knowledge from interaction images and effectively integrates it with both point cloud and image representations to jointly ground object 3D affordance. Specifically, we first devise a Multi-Head Affordance Chain of Thought (MHACoT) reasoning strategy to infer implied invariant geometries and underlying interaction intentions from fine-tuning MLLM, then GREAT employs attention mechanisms to model the correlation between these primitives to form the affordance knowledge dictionary. Following this, we design the Cross-Modal Adaptive Fusion Module (CMAFM) to integrate knowledge into the point cloud features and directly fuse it with image features, leveraging the combined representation to accurately ground 3D object affordance.
@@
Moreover, to further unleash the model’s capability, we extend the Point Image Affordance Dataset (PIAD) [56] to PIADv2, including 24 common affordances, 43 different object categories, over 15K interaction images from diverse scenes and 38K 3D objects with affordance annotations, triple interaction images and over 5 times 3D instances compared to the PIAD (Tab. 1). With its comprehensive scope, PIADv2 offers a robust and reliable testbed for 3D visual affordance.
@@ 2 Related Work
Affordance Grounding. Affordance grounding aims to locate the region of “action possibilities”, which is a link between robot perception and manipulation. Some works ground the object affordance from the 2D data e.g. images and videos [35, 33, 22, 61], while some works leverage natural language understanding to ground affordance regions in 2D data [32, 23, 3]. However, robot manipulation usually requires 3D information of objects, and the 2D affordance grounding obtained from the above works make it difficult to infer the interaction position of 3D objects. With the presentation of several 3D object datasetset [5, 39, 46], some works focus on the 3D object affordance grounding [6, 40, 39], which map semantic affordance to 3D object structure and fail to handle the open-vocabulary scenario.
Open-Vocabulary 3D Affordance Grouding. OVAG presents a substantial chanllenge, aiming to locate object affordance region in arbitrary scenario. Recently, some methods explore the possibility of OVAG. IAGNet [56] utilizes the 2D interaction semantics to guide the grounding of 3D object affordance. LASO [27] employs textual-conditioned affordance queries to isolate afforded segments and injects text clues to point features. OpenAD [42] and OpenKD [50] propose a text-point correlation method for affordance synonym substitutions by utilizing the power of large language encoder clip. Although above methods have made remarkable progress, they are still vulnerable to a limited training semantic space due to the presence of critical learnable parts in the framework. GREAT mitigates this limitation by leveraging geometry-intention collaborative inference with CoT to ground object affordance.
Chain of Thought Prompting with MLLMs. Chain of Thought (CoT) [51, 62] and its variants [59, 60, 21] are proposed to enhance the reasoning capabilities of Multi-modal Large Language Models (MLLMs), which guide the model through multiple logical steps. With the rapid development of MLLMs [4, 16, 65], vision-related [55, 43, 53] methods have made significant progress in collaborating CoT and MLLMs to get the desired results. Some methods explore object affordance at task driven object detection [48], robot manipulation [26] and 2D-level object detection [3], while these methods only describe the egocentric images and then reason through the text or image encoding results as inputs, that can only obtain limited and static knowledge. However, despite the considerable achievements of CoT enhanced MLLMs reasoning, it is challenging to reason about 3D object affordance from interaction images due to the complex and dynamic properties of affordance. To bridge this gap, we fine-tune the InternVL [4] and directly input interaction images with prompts to reveal the geometric attributes and underlying interaction intention.
@@ 3.1 Overview
Given the inputs $\{P,I\}$, where $P\in\mathbb{R}^{N\times 4}$ denotes a point cloud of the object comprising the coordinate ${P_{c}}\in\mathbb{R}^{N\times 3}$ and the affordance annotation ${P_{label}}\in\mathbb{R}^{N\times 1}$, $I\in\mathbb{R}^{3\times H\times W}$ denotes an image. $N$ is the number of points, $H,W$ are the height and width of an image. The goal is to optimize the model $f_{\theta}$ that outputs 3D object affordance $\phi$, expressed as $\phi=f_{\theta}(P_{c},I)$. As shown in Fig. 2, initially, the inputs are sent to ResNet [9] and PointNet++ [44], obtain specific features $\mathbf{F}_{i}\in\mathbb{R}^{C\times H_{1}\times W_{1}},\mathbf{F}_{p}\in\mathbb{R}^{C\times N_{p}}$, and reshape $\mathbf{F}_{i}$ to $\mathbb{R}^{C\times N_{i}}$ ($N_{i}=H_{1}\times W_{1}$). Then, GREAT captures object structure attributes and affordance interaction procedure by fine-tuning MLLM [4] with the Multi-Head Affordance Chain of Thought to reason about interaction images. Next, the features encoded by Roberta [30] are fused through cross-attention mechanism, calculating the $\bar{\mathbf{T}}_{o}$ and interntion feature $\bar{\mathbf{T}}_{a}$ (Sec. 3.2). Afterwards, with $\bar{\mathbf{T}}_{o}$, $\bar{\mathbf{T}}_{a}$ as knowledge dictionaries, GREAT leverages Cross-Modal Adaptive Fusion Module to inject knowledge clues into point features and directly fuse knowledge within image features to obtain fusion features $\mathbf{F}_{tp},\mathbf{F}_{ti}$ (Sec. 3.3). Eventually, $\mathbf{F}_{tp}$ and $\mathbf{F}_{ti}$ are sent to the decoder to ground 3D object affordance $\phi$, the whole process is optimized by a combined loss (Sec. 3.4).
@@
3.2 Multi-Head Affordance Chain of Thought
@@
Object-Head Reasoning for Geometry. It consists of Object Interaction Perception and Geometric Structure Reasoning. First, the model needs to focus on understanding the interactive components of an object within the image, refining its perception of the object’s key parts rather than the entire object. As illustrated in Fig. 2 (a), we design the first prompt as “Point out which part of the object in the image interacts with the person.” Next, since different regions of an object can often perform the same interaction based on shared geometric attributes, the model needs to reason about these features. This allows it to move beyond the constraints of object categories and focus more on the relationship between structure and affordance. Thus, we design the second prompt as “Explain why this part can interact from the geometric structure of the object.
@@
Knowledge Encoding and Integration. For each interaction image, we concatenate the textual descriptions of the interaction components and the object’s geometric attributes inferred from the object-head into a textual sequence. Additionally, we combine the actual interaction and two potential interactions inferred from the affordance-head. Use text encoder Roberta [30] to obtain object geometric knowledge feature $\mathbf{T}_{o}\in\mathbb{R}^{N_{o}\times C}$ and affordance intention knowledge feature $\mathbf{T}_{a}\in\mathbb{R}^{N_{a}\times C}$, where $N_{o},N_{a}$ denote as the number of interaction objects and the number of interaction ways. By applying the cross-attention layer $f_{m}$ to correlate information from the two knowledge repositories and the self-attention layer $f_{\delta}$ to enrich the contextual information, the process aligns the object’s geometric structure attributes with potential interaction intentions, as formulated below:
@@
Figure 3: PIADv2 Dataset. (a) Extensive data examples from PIADv2, the red region in point clouds is the affordance annotation. (b) Category distribution in PIADv2. (c) Confusion matrix between affordance and object categories, where the horizontal axis represents object category and the vertical axis represents affordance category. (d) Ratio of images and point clouds in each affordance category.
@@ 3.3 Cross-Modal Adaptive Fusion Module
Specifically, CMAFM re-represents $\mathbf{F}_{p}\in\mathbb{R}^{C\times N_{p}}$ and $\bar{\mathbf{T}}_{o}\in\mathbb{R}^{N_{o}\times C}$ in the same feature space to further align the local features. $\mathbf{F}_{p}$ is projected to form the query $\mathbf{Q}=\mathbf{F}_{p}\mathbf{W}_{1}$, $\bar{\mathbf{T}}_{o}$ is projected to form the key $\mathbf{K}=\bar{\mathbf{T}}_{o}\mathbf{W}_{2}$ and value $\mathbf{V}=\bar{\mathbf{T}}_{o}\mathbf{W}_{3}$, where $\mathbf{W}_{1\sim 3}$ are projection weights. Next, the cross-attention mechanism integrates these features to extract the interaction context information, formulated as:
@@ 3.4 Decoder and Loss Functions
$\small\mathbf{F}_{\alpha}=f[\Gamma(\bar{\mathbf{F}}_{ti}),\mathbf{F}_{tp}],\phi=\sigma(f_{\phi}(\mathbf{F}_{\alpha})),$ (5)
where $\Gamma$ denotes reshape $\bar{\mathbf{F}}_{ti}$ to $\mathbb{R}^{C\times N}$, $f_{\phi}$ denotes an output head, $\sigma$ denotes the sigmoid function, $\mathbf{F}_{\alpha}\in\mathbb{R}^{C\times N}$ is affordance feature representation and $\phi$ represents the 3D object affordance.

### v2

@@ Abstract
Open-Vocabulary 3D object affordance grounding aims to anticipate “action possibilities” regions on 3D objects with arbitrary instructions, which is crucial for robots to generically perceive real scenarios and respond to operational changes. Existing methods focus on combining images or languages that depict interactions with 3D geometries to introduce external interaction priors. However, they are still vulnerable to a limited semantic space by failing to leverage implied invariant geometries and potential interaction intentions. Normally, humans address complex tasks through multi-step reasoning and respond to diverse situations by leveraging associative and analogical thinking. In light of this, we propose GREAT (GeometRy-intEntion collAboraTive inference) for Open-Vocabulary 3D Object Affordance Grounding, a novel framework that mines the object invariant geometry attributes and performs analogically reason in potential interaction scenarios to form affordance knowledge, fully combining the knowledge with both geometries and visual contents to ground 3D object affordance. Besides, we introduce the Point Image Affordance Dataset v2 (PIADv2), the largest 3D object affordance dataset at present to support the task. Extensive experiments demonstrate the effectiveness and superiority of GREAT. The code and dataset are available at https://yawen-shao.github.io/GREAT/.
@@ 1 Introduction
Figure 1: Difference and Motivation. (a) Object affordance grounding on seen setting. (b) Open-Vocabulary Affordance Grounding (OVAG) with previous paradigms. (c) Observing interaction images, people engage in brainstorming through memory representations, drawing on prior interaction experiences to perform analogical reasoning and infer appropriate actions. (d) OVAG with our geometry-intention collaborative inference with chain-of-thought, step-by-step identifies the interaction part, extracts geometric attributes, reasons about corresponding interaction and brainstorms underlying interaction intentions, jointly grounding the 3D object affordance.
@@
Open-Vocabulary 3D object affordance grounding aims to locate “action possibilities” on objects [14, 8], both for seen and unseen scenarios, identifying specific regions on objects that support certain interactions. This bridges visual perception and physical manipulation for embodied agents and possesses bountiful application scenarios, e.g. robot manipulation [12, 19, 40], scene understanding [27, 11], action anticipation [66, 67, 35] and imitation learning [23, 13].
@@
Recently, most existing methods [6, 17, 2] establish explicit mappings between semantic affordance categories and geometric structures, restricted to predefined seen categories and fail to ground object affordance out of the training categories. Thus, some studies [56, 25, 41, 50, 58, 64] explore grounding object affordance through additional instructions, encompassing combining images or languages that depict interactions with 3D geometries to introduce external interaction priors, and mitigate the generalization gap lead by affordance diversity. Despite their remarkable progress, they are still vulnerable to a limited semantic space by failing to leverage implied invariant geometries among objects with the same affordance, as well as potential correlations among distinct interactions of the same object. As shown in Fig. 1 (a), current paradigms ground the 3D object affordance by aligning the object’s geometric features and instruction modalities, working well in the seen partition. However, such a paradigm relies excessively on the data fed into the model, when testing unseen affordance like “pour” that does not appear in the training set, the model endeavors to categorize it as the seen category “grasp”, shown in Fig. 1 (b).
@@
In cognitive science, studies [7, 16, 62] have shown that humans tackle complex tasks through multi-step reasoning and respond to diverse situations by employing associative and analogical thinking. As demonstrated in Fig. 1 (c), when observing a water pouring scenario, humans employ multi-step reasoning to infer multiple potential interaction procedures and identify geometric properties of objects, indexing relevant knowledge from their brains. Analogous to this procedure, we leverage Multi-modal Large Language Models (MLLMs) [4, 22, 36, 49] to simulate prior knowledge, encompassing visual, linguistic, and other modalities, lifting their superior reasoning and generalization abilities to visual tasks [24, 42]. However, the dynamic and diverse nature of affordances makes it challenging to solely derive complex reasoning outcomes from MLLMs, to address this, a Chain-of-Thought (CoT) inference strategy that mirrors human reasoning processes is designed (Fig. 1 (d)). The strategy step-by-step identifies the interaction primitives, extracts geometric attributes, and reasons about the corresponding interaction actions and intentions, which enables the model to analyze interaction images from multiple perspectives, thereby eliminating the hallucination and ambiguities raised by MLLMs when reasoning interactions.
@@
To achieve this, we present the GREAT, a novel framework that excavates implied affordance knowledge from interaction images and effectively integrates it with both point cloud and image representations to jointly ground object 3D affordance. Specifically, we first devise a Multi-Head Affordance Chain-of-Thought (MHACoT) reasoning strategy to infer implied invariant geometries and underlying interaction intentions from fine-tuning MLLM, then GREAT employs attention mechanisms to model the correlation between these primitives to form the affordance knowledge dictionary. Following this, we design the Cross-Modal Adaptive Fusion Module (CMAFM) to integrate knowledge into the point cloud features and directly fuse it with image features, leveraging the combined representation to accurately ground 3D object affordance.
@@
Moreover, to further unleash the model’s capability, we extend the Point Image Affordance Dataset (PIAD) [56] to PIADv2, including 24 common affordances, 43 different object categories, over 15K interaction images from diverse scenes and 38K 3D objects with affordance annotations, triple interaction images and over 5 times 3D instances compared to the PIAD (Tab. 1).
@@ 2 Related Work
Affordance Grounding. Affordance grounding aims to locate the region of “action possibilities”, which is a link between robot perception and manipulation. Some works ground the object affordance from the 2D data e.g. images and videos [33, 31, 63, 61, 34, 21], while some works leverage natural language understanding to ground affordance regions in 2D data [30, 21, 3, 44]. However, robot manipulation usually requires 3D information of objects, and the 2D affordance grounding obtained from the above works make it difficult to infer the interaction position of 3D objects. With the presentation of several 3D object datasets [5, 38, 46], some works focus on the 3D object affordance grounding [6, 39, 38], which map semantic affordance to 3D object structure and fail to handle the open-vocabulary scenario.
Open-Vocabulary 3D Affordance Grouding. OVAG presents a substantial challenge, aiming to locate object affordance region in arbitrary scenario. Recently, some methods explore the possibility of OVAG. IAGNet [56] utilizes the 2D interaction semantics to guide the grounding of 3D object affordance. LASO [25] employs textual-conditioned affordance queries to isolate afforded segments and injects text clues to point features. OpenAD [41] and OpenKD [50] propose a text-point correlation method for affordance synonym substitutions by utilizing the power of large language encoder clip. Although above methods have made remarkable progress, they are still vulnerable to a limited training semantic space due to the presence of critical learnable parts in the framework. GREAT mitigates this limitation by leveraging geometry-intention collaborative inference with CoT to ground object affordance.
Chain-of-Thought Prompting with MLLMs. Chain-of-Thought (CoT) [51, 65] and its variants [59, 60, 20] are proposed to enhance the reasoning capabilities of Multi-modal Large Language Models (MLLMs), which guide the model through multiple logical steps. With the rapid development of MLLMs [4, 15, 68], vision-related [55, 42, 53] methods have made significant progress in collaborating CoT and MLLMs to get the desired results. Some methods explore object affordance at task driven object detection [48], robot manipulation [24] and 2D-level object detection [3], while these methods only describe the egocentric images and then reason through the text or image encoding results as inputs, that can only obtain limited and static knowledge. However, despite the considerable achievements of CoT enhanced MLLMs reasoning, it is challenging to reason about 3D object affordance from interaction images due to the complex and dynamic properties of affordance. To bridge this gap, we fine-tune the InternVL [4] and directly input interaction images with prompts to reveal the geometric attributes and underlying interaction intentions.
@@ 3.1 Overview
Given the inputs $\{P,I\}$, where $P\in\mathbb{R}^{N\times 4}$ denotes a point cloud of the object comprising the coordinate ${P_{c}}\in\mathbb{R}^{N\times 3}$ and the affordance annotation ${P_{label}}\in\mathbb{R}^{N\times 1}$, $I\in\mathbb{R}^{3\times H\times W}$ denotes an image. $N$ is the number of points, $H,W$ are the height and width of an image. The goal is to optimize the model $f_{\theta}$ that outputs 3D object affordance $\phi$, expressed as $\phi=f_{\theta}(P_{c},I)$. As shown in Fig. 2, initially, the inputs are sent to ResNet [9] and PointNet++ [43], obtain specific features $\mathbf{F}_{i}\in\mathbb{R}^{C\times H_{1}\times W_{1}},\mathbf{F}_{p}\in\mathbb{R}^{C\times N_{p}}$, and reshape $\mathbf{F}_{i}$ to $\mathbb{R}^{C\times N_{i}}$ ($N_{i}=H_{1}\times W_{1}$). Then, GREAT captures object structure attributes and affordance interaction procedure by fine-tuning MLLM [4] with the multi-head affordance chain-of-thought to reason about interaction images. Next, the features encoded by Roberta [28] are fused through cross-attention mechanism, calculating object geometric feature $\bar{\mathbf{T}}_{o}$ and affordance intention feature $\bar{\mathbf{T}}_{a}$ (Sec. 3.2). Afterwards, with $\bar{\mathbf{T}}_{o}$, $\bar{\mathbf{T}}_{a}$ as knowledge dictionaries, GREAT leverages Cross-Modal Adaptive Fusion Module to inject knowledge clues into point features and directly fuse knowledge within image features to obtain fusion features $\mathbf{F}_{tp},\mathbf{F}_{ti}$ (Sec. 3.3). Eventually, $\mathbf{F}_{tp}$ and $\mathbf{F}_{ti}$ are sent to the decoder to ground 3D object affordance $\phi$, the whole process is optimized by a combined loss (Sec. 3.4).
@@
3.2 Multi-Head Affordance Chain-of-Thought
@@
Object-Head Reasoning for Geometry. It consists of Object Interaction Perception and Geometric Structure Reasoning. First, the model needs to focus on understanding the interactive components of an object within the image, refining its perception of the object’s key parts rather than the entire object. As illustrated in Fig. 2 (a), we design the first prompt as “Point out which part of the object in the image interacts with the person.” Next, since similar regions of different objects can often perform the same interaction based on shared geometric attributes, the model needs to reason about these features. This allows it to move beyond the constraints of object categories and focus more on the relationship between structure and affordance. Thus, we design the second prompt as “Explain why this part can interact from the geometric structure of the object.
@@
Knowledge Encoding and Integration. For each interaction image, we concatenate the textual descriptions of the interaction components and the object’s geometric attributes inferred from the object-head into a textual sequence. Additionally, we combine the actual interaction and two potential interactions inferred from the affordance-head. The text encoder Roberta [28] is used to obtain object geometric knowledge feature $\mathbf{T}_{o}\in\mathbb{R}^{N_{o}\times C}$ and affordance intention knowledge feature $\mathbf{T}_{a}\in\mathbb{R}^{N_{a}\times C}$, where $N_{o},N_{a}$ denote as the number of interaction objects and the number of interaction ways. By applying the cross-attention layer $f_{m}$ to correlate information from the two knowledge repositories and the self-attention layer $f_{\delta}$ to enrich the contextual information, the process aligns the object’s geometric structure attributes with potential interaction intentions, as formulated below:
@@
Figure 3: PIADv2 Dataset. (a) Extensive data examples from PIADv2, the red region in point clouds is the affordance annotation. (b) Category distribution in PIADv2. (c) Confusion matrix between affordance and object categories, where the horizontal axis represents object category and the vertical axis represents affordance category. (d) Ratio of images and point clouds in each object category.
@@ 3.3 Cross-Modal Adaptive Fusion Module
Specifically, CMAFM re-represents $\mathbf{F}_{p}\in\mathbb{R}^{C\times N_{p}}$ and $\bar{\mathbf{T}}_{o}\in\mathbb{R}^{N_{o}\times C}$ in the same feature space to further align the local features. $\mathbf{F}_{p}$ is projected to form the query $\mathbf{Q}=\mathbf{F}_{p}\mathbf{W}_{1}$, $\bar{\mathbf{T}}_{o}$ is projected to form the key $\mathbf{K}=\bar{\mathbf{T}}_{o}\mathbf{W}_{2}$ and value $\mathbf{V}=\bar{\mathbf{T}}_{o}\mathbf{W}_{3}$, where $\mathbf{W}_{1\sim 3}$ are projection weights. The cross-attention mechanism integrates these features to extract the interaction context information, formulated as:
@@ 3.4 Decoder and Loss Functions
$\small\mathbf{F}_{\alpha}=f[\Gamma({\mathbf{F}}_{ti}),\mathbf{F}_{tp}],\phi=\sigma(f_{\phi}(\mathbf{F}_{\alpha})),$ (5)
where $\Gamma$ denotes reshape ${\mathbf{F}}_{ti}$ to $\mathbb{R}^{C\times N}$, $f_{\phi}$ denotes an output head, $\sigma$ denotes the sigmoid function, $\mathbf{F}_{\alpha}\in\mathbb{R}^{C\times N}$ is affordance feature representation and $\phi\in\mathbb{R}^{N\times 1}$ represents the 3D object affordance.